OpenStudy (anonymous):

If the first term of an arithmetic sequence is 10 and the 13th term is 58, what is the common difference?

OpenStudy (anonymous):

first term=10=a (let) 13th term=58 now in arimetic sequence n th term is=a+(n-1)d d=common difference so, 13th term=a+(13-1)d=a+12d now, a+12d=58 =>10+12d=58 =>d=4 so,common difference is 4.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome...